I asked this question before but I didn't really understand what was the answers for EACH of the blanks listed in my question, so I am asking it again. A 1.24 M solution of KI has a density of 1.15 g/cm3. Assume 1.00 liter of 1.24 M KI solution. The mass of 1.00 L of solution is ______grams. The number of grams of KI in the 1 liter is _____ grams. The number of grams of water in the 1 L is _____grams. The molality of the KI solution is _____ molal. The molal freezing point depression constant for water in _____deg C kg/molal. The van't Hoff factor for complete dissociation is______ . For complete dissociation of KI the freezing point depression is _____deg C. If the actual freezing point depression is -4.46 deg C, the van't Hoff factor value is _____and the percent of dissociation is _____ (2 S.F.). Molality is moles of solute per _____ of solvent.
Answer #1

Given

Molarity = 1.24 M = 1.24 mol/L of KI

density = 1.15 g/cm3 = 1.15 g/ml

Volume = 1L = 1000 ml of solution

mass = Volume * density = 1000 ml * 1.15 g/ml = 1150 g Answer 1

No. of moles of KI = Molarity of KI * Volume of solution = 1.24 mol/L * 1 L = 1.24 moles

Molar mass of KI = 166 g/mol

Mass of KI = No. of moles * Molar mass = 1.24 moles * 166 g/mol = 205.84 g Answer 2

No. of grams of water = mass of solution - mass of KI = 1150 g - 205.84 g = 944.16 g Answer 3

Molality = No. of moles of solute / Mass of solvent = 1.24 moles / 0.94416 kg = 1.31 mol/Kg ( or m) Answer 4

Molal freezing point depression for water Kb= 1.86 C / m Answer 5

van't hoff factor i = 1 Answer 6

for complete diassocaiton KI freezing point depression

\DeltaT = Kb * m * i = 1.86 C /m * 1.31 m * 1 = - 2.4 C Answer 7

negative sign is because freezing point will always reduce

if\DeltaT is - 4.46

ratio is 4.46 / 2.4 = 1.86

so all others are remaining same

Vant hoff factor will be 1.86 Answer 8

KI ----> K+ + I-

1 moles 1 mol 1mol

after diassociation there should be two moles

percent diassociation is (1.86 /2 ) * 100 % = 93 % Answer 9

Molality is moles of solute per Kg of solvent Answer 10